OpenStudy (anonymous):

simplify using only positive exponents in the final answer...im pretty sure i have this right but i thought i'd check it... 12a^3bc^4/18a^5b^2c

OpenStudy (dumbcow):

what did you come up with?

OpenStudy (dumbcow):

ok if you reduced 12/18 to 2/3 and subtracted the exponents then you would be correct

OpenStudy (dumbcow):

so.. what did you end up with on the numerator

OpenStudy (dumbcow):

are you there?

OpenStudy (dumbcow):

nothing huh well here's the solution \[\frac{12a^{3}bc^{4}}{18a^{5}b^{2}c} = \frac{2c^{4-1}}{3a^{5-3}b^{2-1}} = \frac{2c^{3}}{3a^{2}b}\]

OpenStudy (anonymous):

ahh im sorry i was getting dinner haha. and that's what i had!! thank you so much :)

OpenStudy (dumbcow):

no prob, you never know if they are really there or not, all you see is there icon awesome, good job to you then :)